OpenStudy (anonymous):

Given cosθ=–15 and tan θ > 0, find sin θ. 45 −45 26√5 −26√5

Parth (parthkohli):

@satellite73 Here's another outrageous question.

Parth (parthkohli):

I hope teachers learn that \(-1 \le \cos\theta \le 1.\)